OpenStudy (anonymous):

Solve the equation on the interval[0,2pi] sin^2x+sin x=0

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (anonymous):

assume sin x be t then find the soln

OpenStudy (anonymous):

How do I factor?

OpenStudy (tkhunny):

Why not just factor? y^2 + y = 0 y(y+1) = 0 Frog^2 + Frog = 0 Frog(Frog + 1) = 0 sin^2(x) + sin(x) = 0 sin(x)(sin(x)+1) = 0 You should remember your algebra! No good to forget!